# German General Personas: A Survey-Derived Persona Prompt Collection for Population-Aligned LLM Studies

### v1

@@ 4.1.  Experimental Setup
Possible use cases and tasks of representative persona collections are plentiful. In this paper we focus on evaluating the utility of GGP for simulating the response distribution of survey participants across a diverse set of outcome variables.
@@
This choice of baseline puts the persona-based zero-shot LLM approach at a disadvantage. While our approach always acts on a single persona description as input, the random forest classifiers are fitted on training samples representing persona descriptions and their corresponding responses for the respective prediction tasks, with training dataset sizes systematically ranging from $n=2$ to $n=2,048$. We reserve 2020% of respondents as fixed test set, and randomly select $n$ of the remaining respondents for training the baseline classifiers. The setting of classifier hyperparameters follows miranda2025simulating and is reported in Appendix C. We compare the classifier predictions with the groundtruth response distribution of the test set respondents. The LLMs do not have access to any training samples.
@@
We measure the performance of our approach as well as of the different baselines by calculating the Jensen-Shannon Distance (JSD) between the predicted and the actual response distribution of each sampled outcome variable. The predicted response distribution is created by aggregating the model predictions for a given variable across all participants, and the actual response distribution is the aggregate of the participants’ responses found in the ALLBUS. This procedure is in line with the suggestion offered by sorensen2024position for the evaluation of distributionally pluralistic models.
@@
As discussed above, the GGP is representative of the German population because it mirrors the randomly sampled participants of the ALLBUS.
In the experiments in which we are interested in the effects of representativity on the ability of persona-prompted LLMs to replicate the population’s response distribution, we systematically create persona collections that are by design unrepresentative, e.g., oversampled on specific attributes. We sample $500$ participants from the highest income class to oversample on income, $500$ participants that lean conservative (based on their responses to the TOP-$2$ attributes, see Figure 1) to oversample on ideological leaning, and $500$ students to mimic a typical convenience sample population in social science studies sears1986college.
In addition, we run the same survey response tasks with the popular, but empirically not grounded PersonaHub collection ge2024scaling, as well as using no personas at all. To ensure comparability of the results, we randomly sampled a subset of $500$ personas from the representative GGP. By including multiple baselines we try to identify whether GGP’s representativeness offers any advantages over non-representative alternatives.
@@ 4.2.  Results
We show that persona-prompting with the GGP offers advantages over traditional statistical methods for survey response predictions in situations marked by data sparsity with few survey participants and little information about them available (Figure 2), and that increasing the level of information reflected in persona descriptions does not necessarily lead to improved performance (Figure 3).
Additionally, we show that using the GGP leads to predicted response distributions that are better aligned with the survey response distribution than those produced using established baselines in $13$ out of $27$ tasks and across five out of nine topics (Figure 4). Finally, we show that the representativity of the GGP seems to have only little influence on the level of alignment (Figure 5).
@@
Figure 4: Alignment comparison across different topics. We compare the Jensen-Shannon Distance (JSD) between the survey response distribution and the response distribution generated using the best GGP configuration as well as the response distributions from random forest baselines with different training set sizes. Using the GGP and thus no input other than a single persona description produces response distributions that are better aligned with survey responses than the best random forest baselines across five out of nine different topics.
@@
We examine this observation more closely in Figure 3, which shows that there is no monotonous relationship between the number of TOP-$k$ attributes used for constructing the personas and the resulting alignment between the generated and the actual response distributions. Surprisingly, it is again the version of the GGP that uses only the TOP-$2$ persona attributes that leads to the best alignment across all constellations we tested. For the best-performing LLM, we even see a clear tendency for the alignment to degrade with higher numbers of included persona attributes, highlighting a potential signal-noise trade-off when including more and increasingly unimportant persona attributes. Figure 3 also reveals the general tendency that the best possible alignment improves with the size of the persona-prompted LLM used for generating the synthetic response distribution. Only the unlikely strong alignment of the relatively small Llama-3.1-8B-Instruct prompted with only socio-demographic information breaks this trend, warranting closer investigation.
Figure 5: Alignment comparison with unrepresentative persona collections. We compare the average Jensen-Shannon Distance of response distributions generated with representative (GGP), unrepresentative (baselines), and no persona collections to the survey response distribution. All synthetic responses are generated using Llama-3.3-70B-Instruct. The unrepresentative baselines and the GGP cluster closely together in terms of their JSD scores, indicating that representativity only has a small influence on the alignment.
@@
In Figure 4, we compare the performance of persona-prompting LLMs using the GGP against random forest classifiers fitted on increasing numbers of training samples, now showing the disaggregated Jensen-Shannon Distances measured across nine different topics. Based on Figure 3, we use the best performing constellation for predicting response distributions, i.e., Llama-3.3-70B-Instruct prompted with TOP-$2$ personas, and increase the number of training samples for the baseline classifiers up to $n=512$, after which we do not observe any more performance improvements. We report the average JSD between the predicted and actual survey response distributions across the three outcome variables randomly sampled for each topic.
Our persona-prompted LLM, receiving only the TOP-$2$ persona attributes in the prompt, outperforms even the best informed baseline with access to up to $n=512$ of training samples and the same set of attributes in predicting the response distribution for $13$ out of $27$ different outcome variables. On the topic-level, the LLM-based approach turns out to be particularly strong in predicting response distributions in the categories Social Capital, Economic Situation, Religion, Value & Life Goals, and Ethnocentrism, where it achieves lower average JSD to the actual response distribution than even the best random forest classifiers. The absolute alignment using the GGP to persona-prompt LLMs is best for the topics Social Capital and Economic Situation, and worst for the topics Social Inequality and Morality.
@@
We run the same experiments using unrepresentative—oversampled on specific attributes and not empirically grounded—persona collections as baselines, as described in Section 4.1. In this comparison, there is no clear pattern showing that response distributions predicted using the representative GGP are consistently better aligned than the baselines they are compared against. Instead, Figure 5 shows that the different persona collections are generally scattered closely together around similar alignment scores across different topics, indicating that the predictive performance is highly topic dependent.
For the topics Social Capital, Values & Life Goals, Political Tendency and Economic Situation, the persona-prompted response distributions are closer aligned with the survey response distribution than the response distributions generated using no personas, indicating that the TOP-$2$ attributes used in the persona descriptions might have been helpful in correctly predicting the responses in these areas. The baseline oversampled on conservatives (as derived from the TOP-$2$ attributes) deviates most strongly from the GGP, strengthening our assumption that these attributes have the highest impact on the prediction of other ALLBUS variables. The fact that the empirically ungrounded PersonaHub collection aligns so closely with the different persona collections based on the ALLBUS is surprising.
@@ 5.  Discussion
This work introduces German General Personas - a novel prompt collection designed to evaluate and enhance the use of persona-based prompting for modeling response distributions of populations. Our preliminary experiments demonstrate the current ability of persona-based approaches to outperform traditional response prediction and imputation methods, particularly in situations characterized by limited data availability. These findings suggest that persona prompting can serve as a viable and flexible alternative in low-resource contexts where conventional supervised methods may struggle.
A key consideration in developing German General Personas is achieving a balance between generalizability—ensuring that the collection remains applicable across a wide range of downstream tasks—and specificity, which is essential for strong performance in domain-specific applications. In this work, we propose leveraging general population surveys to construct empirically informed persona collections. Interestingly, our results indicate that personas based on a small subset of key variables outperform those that incorporate all available survey variables. This suggests that including only the most informative attributes may enhance both efficiency and predictive accuracy. In our preliminary analysis we only considered the average importance of variables for all other outcome variable. An important direction for future research is to explore how persona descriptions can be systematically tuned to optimally align with observed response patterns.
Recent work has begun fine-tuning large language models (LLMs) not only to generate single “correct” answers, but to minimize the divergence between the model’s predicted response distributions and empirical human survey data from sources such as the World Values Survey and Pew Global Attitudes Survey (cao2025specializing; suh2025language). However, these supervised distribution-fitting methods require substantial computational resources and access to proprietary model weights, making them less applicable in low-resource or restricted-access settings. In contrast, persona prompting offers a more lightweight and adaptable alternative that can be applied even to publicly available models. The performance of persona-based methods can potentially further be improved through few-shot approaches and with the release of increasingly capable language models. This highlights a promising avenue for future work, combining the interpretability and flexibility of prompting with the precision gains typically associated with fine-tuned models.
German General Personas can act as a valuable resource that supports systematic comparisons of population-alignment of LLMs within different tasks (e.g., survey response generation, discussion simulations or behavior predictions).
@@ 6.  Limitations
The idea of turning general social surveys such as the ALLBUS into empirically grounded persona collection comes with a number of limitations. First, surveys only offer a snapshot of a population that dynamically changes over time, both in its composition (e.g., due to migration or demographic change) as well as in the attitudes and beliefs held by individuals. To explicitly mark this temporal dependency, we date the released GGP through versioning. A partial alleviation of this limitation could also be in the use of reweighting techniques, which we leave for future work to explore. Second, while we rule out the possibility that any of our tested LLMs has had access to the ALLBUS data we use, another (indirect) form of data leakage could occur across different releases of the ALLBUS, given that such surveys to a large degree run the same questions in every iteration. Future work could thus investigate performance differences between GGP versions created from and evaluated on current and past releases of the ALLBUS. Third, missing information at an individual level introduces a trade-off between the representativity of the persona collection and the completeness of the individual persona descriptions. In our experiments, we prioritize representativity by including all personas instead of selecting only those that are complete, i.e., have information on all attributes available. By releasing the GGP with all personas, including those that are incomplete, we allow users of the collection to resolve this trade-off as they see fit—they could prioritize representativity by working with the collection as is, or they could prioritize persona completeness by selecting only personas with full information. Future work could explore how imputation procedures, possibly LLM-based castricato2025persona, might help to resolve this trade-off. Fourth, our attribute selection procedure does not account for correlations and possible multicollinearity between predictors. Future work could test methods that evaluate all possible subsets of $k$ attributes to improve upon the sets of TOP-$k$ attributes currently used for persona creation. Lastly, we are relying exclusively on survey data, both for the creation of persona descriptions and as a source of human groundtruth to evaluate our approach against. However, there is some evidence that survey data itself might be limited in accurately capturing human attitudes and behaviors. Future work could on the one hand explore the inclusion of behavioral data in persona prompts as a more immediate proxy of human behavior, and on the other hand search additional sources of groundtruth (such as election outcomes) to evaluate against.
@@
Appendix A German General Personas
@@
Figure 6: Grounding the German General Personas (GGP) in the ALLBUS survey. We construct individual persona prompts for each ALLBUS participant, varying size and composition via available attributes. A global variable importance ranking informs the selection of the $k$ most important attributes (TOP-$k$). Personas comprise a fixed block of core socio-demographics and a more extensible block of TOP-$k$ attributes that allow varying information content. The GGP is available in JSON- and full-text formats.
@@ B.1.  Experimental Setup
The provided German persona prompt is an example of the sociodemographic attributes contained in each persona as well as the two identified most relevant attributes in the ALLBUS,
@@
instructs the LLM to adopt a specific persona. This persona is a 50-year-old male from West Germany with an intermediate school leaving certificate, living in a large city and working in assembly with a net monthly income of 1250 - 1374 Euros. He supports limiting immigration from non-EU countries and is willing to voice his political opinions. The prompt then asks for this persona’s reaction to a question about the responsibility of having children (variable name: lp04), with the options being "agree" or "disagree".
The English translation is:
@@
Table 3 shows the rates with which the different LLMs used for generating synthetic response distributions failed to generate valid responses.
@@
We carried out the experiments of predicting attributes with RPPC’s synthetic responses on a high-performance computing cluster and a local server equipped with NVIDIA H100 (80GB) GPUs as well as Blackwell GPUs. The total runtime on the computing cluster for **********. To accommodate larger models on available hardware, we applied 8-bit quantization to Llama-3.3-70B-Instruct. Smaller models were run without quantization.
Further, we use vLLM for a more efficient memory management and a higher throughput as this packages enables sufficient batching of many requests at a time kwon2023efficient.
@@ Appendix C Evaluation
C.1.  Hyperparameters for Random Forest Classifier
The Random Forest models utilized for classification tasks throughout all experiments were instantiated from the scikit-learn library’s RandomForestClassifier class. We selected the hyperparameters according to the findings of miranda2025simulating who used the classifier in a similar task. To ensure consistency and comparability across experiments, the following hyperparameters were fixed:

### v2

@@ 4.1.  Experimental Setup
Possible use cases and tasks of representative persona collections are plentiful. In this paper we focus on evaluating the utility of GGSS Personas for simulating the response distribution of survey participants across a diverse set of outcome variables.
@@
This choice of baseline puts the persona-based zero-shot LLM approach at a disadvantage. While our approach always acts on a single persona description as input, the random forest classifiers are fitted on training samples representing persona descriptions and their corresponding responses for the respective prediction tasks, with training dataset sizes systematically ranging from $n=2$ to $n=2,048$. We reserve 2020% of respondents as fixed test set, and randomly select $n$ of the remaining respondents for training the baseline classifiers. The setting of classifier hyperparameters follows Miranda and Balbi (2025) and is reported in Appendix C. We compare the classifier predictions with the groundtruth response distribution of the test set respondents. The LLMs do not have access to any training samples.
@@
We measure the performance of our approach as well as of the different baselines by calculating the Jensen-Shannon Distance (JSD) between the predicted and the actual response distribution of each sampled outcome variable. The predicted response distribution is created by aggregating the model predictions for a given variable across all participants, and the actual response distribution is the aggregate of the participants’ responses found in the ALLBUS. This procedure is in line with the suggestion offered by Sorensen et al. (2024) for the evaluation of distributionally pluralistic models.
@@
As discussed above, the GGSS Personas is representative of the German population because it mirrors the randomly sampled participants of the ALLBUS.
In the experiments in which we are interested in the effects of representativity on the ability of persona-prompted LLMs to replicate the population’s response distribution, we systematically create persona collections that are by design unrepresentative, e.g., oversampled on specific attributes. We sample $500$ participants from the highest income class to oversample on income, $500$ participants that lean conservative (based on their responses to the TOP-$2$ attributes, see Figure 1) to oversample on ideological leaning, and $500$ students to mimic a typical convenience sample population in social science studies Sears (1986).
In addition, we run the same survey response tasks with the popular, but empirically not grounded PersonaHub collection Ge et al. (2024), as well as using no personas at all. To ensure comparability of the results, we randomly sampled a subset of $500$ personas from the representative GGSS Personas. By including multiple baselines we try to identify whether GGSS Personas’s representativeness offers any advantages over non-representative alternatives.
@@ 4.2.  Results
We show that persona-prompting with the GGSS Personas offers advantages over traditional statistical methods for survey response predictions in situations marked by data sparsity with few survey participants and little information about them available (Figure 2), and that increasing the level of information reflected in persona descriptions does not necessarily lead to improved performance (Figure 3).
Additionally, we show that using the GGSS Personas leads to predicted response distributions that are better aligned with the survey response distribution than those produced using established baselines in $13$ out of $27$ tasks and across five out of nine topics (Figure 4). Finally, we show that the representativity of the GGSS Personas seems to have only little influence on the level of alignment (Figure 5).
@@
Figure 4: Alignment comparison across different topics. We compare the Jensen-Shannon Distance (JSD) between the survey response distribution and the response distribution generated using the best GGSS Personas configuration as well as the response distributions from random forest baselines with different training set sizes. Using the GGSS Personas and thus no input other than a single persona description produces response distributions that are better aligned with survey responses than the best random forest baselines across five out of nine different topics.
@@
We examine this observation more closely in Figure 3, which shows that there is no monotonous relationship between the number of TOP-$k$ attributes used for constructing the personas and the resulting alignment between the generated and the actual response distributions. Surprisingly, it is again the version of the GGSS Personas that uses only the TOP-$2$ persona attributes that leads to the best alignment across all constellations we tested. For the best-performing LLM, we even see a clear tendency for the alignment to degrade with higher numbers of included persona attributes, highlighting a potential signal-noise trade-off when including more and increasingly unimportant persona attributes. Figure 3 also reveals the general tendency that the best possible alignment improves with the size of the persona-prompted LLM used for generating the synthetic response distribution. Only the unlikely strong alignment of the relatively small Llama-3.1-8B-Instruct prompted with only socio-demographic information breaks this trend, warranting closer investigation.
Figure 5: Alignment comparison with unrepresentative persona collections. We compare the average Jensen-Shannon Distance of response distributions generated with representative (GGSS Personas), unrepresentative (baselines), and no persona collections to the survey response distribution. All synthetic responses are generated using Llama-3.3-70B-Instruct. The unrepresentative baselines and the GGSS Personas cluster closely together in terms of their JSD scores, indicating that representativity only has a small influence on the alignment.
@@
In Figure 4, we compare the performance of persona-prompting LLMs using the GGSS Personas against random forest classifiers fitted on increasing numbers of training samples, now showing the disaggregated Jensen-Shannon Distances measured across nine different topics. Based on Figure 3, we use the best performing constellation for predicting response distributions, i.e., Llama-3.3-70B-Instruct prompted with TOP-$2$ personas, and increase the number of training samples for the baseline classifiers up to $n=512$, after which we do not observe any more performance improvements. We report the average JSD between the predicted and actual survey response distributions across the three outcome variables randomly sampled for each topic.
Our persona-prompted LLM, receiving only the TOP-$2$ persona attributes in the prompt, outperforms even the best informed baseline with access to up to $n=512$ of training samples and the same set of attributes in predicting the response distribution for $13$ out of $27$ different outcome variables. On the topic-level, the LLM-based approach turns out to be particularly strong in predicting response distributions in the categories Social Capital, Economic Situation, Religion, Value & Life Goals, and Ethnocentrism, where it achieves lower average JSD to the actual response distribution than even the best random forest classifiers. The absolute alignment using the GGSS Personas to persona-prompt LLMs is best for the topics Social Capital and Economic Situation, and worst for the topics Social Inequality and Morality.
@@
We run the same experiments using unrepresentative—oversampled on specific attributes and not empirically grounded—persona collections as baselines, as described in Section 4.1. In this comparison, there is no clear pattern showing that response distributions predicted using the representative GGSS Personas are consistently better aligned than the baselines they are compared against. Instead, Figure 5 shows that the different persona collections are generally scattered closely together around similar alignment scores across different topics, indicating that the predictive performance is highly topic dependent.
For the topics Social Capital, Values & Life Goals, Political Tendency and Economic Situation, the persona-prompted response distributions are closer aligned with the survey response distribution than the response distributions generated using no personas, indicating that the TOP-$2$ attributes used in the persona descriptions might have been helpful in correctly predicting the responses in these areas. The baseline oversampled on conservatives (as derived from the TOP-$2$ attributes) deviates most strongly from the GGSS Personas, strengthening our assumption that these attributes have the highest impact on the prediction of other ALLBUS variables. The fact that the empirically ungrounded PersonaHub collection aligns so closely with the different persona collections based on the ALLBUS is surprising.
@@ 5.  Discussion
This work introduces German General Social Survey Personas - a novel prompt collection designed to evaluate and enhance the use of persona-based prompting for modeling response distributions of populations. Our preliminary experiments demonstrate the current ability of persona-based approaches to outperform traditional response prediction and imputation methods, particularly in situations characterized by limited data availability. These findings suggest that persona prompting can serve as a viable and flexible alternative in low-resource contexts where conventional supervised methods may struggle.
A key consideration in developing German General Social Survey Personas is achieving a balance between generalizability—ensuring that the collection remains applicable across a wide range of downstream tasks—and specificity, which is essential for strong performance in domain-specific applications. In this work, we propose leveraging general population surveys to construct empirically informed persona collections. Interestingly, our results indicate that personas based on a small subset of key variables outperform those that incorporate all available survey variables. This suggests that including only the most informative attributes may enhance both efficiency and predictive accuracy. In our preliminary analysis we only considered the average importance of variables for all other outcome variable. An important direction for future research is to explore how persona descriptions can be systematically tuned to optimally align with observed response patterns.
Recent work has begun fine-tuning large language models (LLMs) not only to generate single “correct” answers, but to minimize the divergence between the model’s predicted response distributions and empirical human survey data from sources such as the World Values Survey and Pew Global Attitudes Survey (Cao et al., 2025; Suh et al., 2025). However, these supervised distribution-fitting methods require substantial computational resources and access to proprietary model weights, making them less applicable in low-resource or restricted-access settings. In contrast, persona prompting offers a more lightweight and adaptable alternative that can be applied even to publicly available models. The performance of persona-based methods can potentially further be improved through few-shot approaches and with the release of increasingly capable language models. This highlights a promising avenue for future work, combining the interpretability and flexibility of prompting with the precision gains typically associated with fine-tuned models.
German General Social Survey Personas can act as a valuable resource that supports systematic comparisons of population-alignment of LLMs within different tasks (e.g., survey response generation, discussion simulations or behavior predictions).
@@ 6.  Limitations
The idea of turning general social surveys such as the ALLBUS into empirically grounded persona collection comes with a number of limitations. First, surveys only offer a snapshot of a population that dynamically changes over time, both in its composition (e.g., due to migration or demographic change) as well as in the attitudes and beliefs held by individuals. To explicitly mark this temporal dependency, we date the released GGSS Personas through versioning. A partial alleviation of this limitation could also be in the use of reweighting techniques, which we leave for future work to explore.
Second, while we rule out the possibility that any of our tested LLMs has had access to the ALLBUS data we use, another (indirect) form of data leakage could occur across different releases of the ALLBUS, given that such surveys to a large degree run the same questions in every iteration. Future work could thus investigate performance differences between GGSS Personas versions created from and evaluated on current and past releases of the ALLBUS.
Third, missing information at an individual level introduces a trade-off between the representativity of the persona collection and the completeness of the individual persona descriptions. In our experiments, we prioritize representativity by including all personas instead of selecting only those that are complete, i.e., have information on all attributes available. By releasing the GGSS Personas with all personas, including those that are incomplete, we allow users of the collection to resolve this trade-off as they see fit—they could prioritize representativity by working with the collection as is, or they could prioritize persona completeness by selecting only personas with full information. Future work could explore how imputation procedures, possibly LLM-based Castricato et al. (2025), might help to resolve this trade-off.
Fourth, our attribute selection procedure does not account for correlations and possible multicollinearity between predictors. Future work could test methods that evaluate all possible subsets of $k$ attributes to improve upon the sets of TOP-$k$ attributes currently used for persona creation.
In addition, LLM-generated closed-ended survey responses are susceptible to minimal perturbations, such as the positioning of the answer option Rupprecht et al. (2025); Tjuatja et al. (2024). Special attention should be put on the robustness of synthetic survey responses by applying prompt perturbations to ensure that the results are not only artifacts of the prompting style.
Lastly, we are relying exclusively on survey data, both for the creation of persona descriptions and as a source of human groundtruth to evaluate our approach against. However, there is some evidence that survey data itself might be limited in accurately capturing human attitudes and behaviors. Future work could on the one hand explore the inclusion of behavioral data in persona prompts as a more immediate proxy of human behavior, and on the other hand search additional sources of groundtruth (such as election outcomes) to evaluate against.
@@
Acknowledgement
We thank all anonymous reviewers for their thoughtful feedback and constructive suggestions, which greatly improved the quality of this work. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
Appendix A German General Social Survey Personas
@@
Figure 6: German Version of Figure 1: Grounding the German General Social Survey Personas (GGSS Personas) in the ALLBUS survey.
@@ B.1.  Experimental Setup
Below, we provide an example of a German persona prompt, featuring the core set of socio-demographic attributes as well as the two attributes identified to be most important (TOP-2) according to our variable importance ranking.
@@
This prompt structure instructs the LLM to take on the perspective of the persona described next, provides the survey question the model is supposed to answer as well as the available response options, and sets the model to up to generate its response in the required format.
In this example, the persona is a 50-year-old male from West Germany with an intermediate school leaving certificate, living in a large city and working in assembly with a net monthly income of 1250 - 1374 Euros, supporting limiting immigration from non-EU countries, and willing to voice his political opinions among friends and at work. The prompt then asks for this persona’s reaction to a question about the responsibility of having children (variable name: lp04), with the options being "agree" or "disagree".
The English translation of the provided prompt is the following:
@@
Decoding Constraints and Invalid Responses
We generated responses from the whole token distribution. We did not apply any decoding constraints such as the Top-P, Top-K sampling or temperature setting. We do acknowledge that these might impact the response generation.
A response is valid when the language model generates only one token that reflects one response option on the answer option scale. Table 3 shows the rates with which the different LLMs used for generating synthetic response distributions failed to generate valid responses.
@@
We carried out the experiments of predicting attributes with GGSS Personas’s synthetic responses on a high-performance computing cluster and a local server equipped with NVIDIA H100 (80GB) GPUs and Blackwell GPUs. To accommodate larger models on available hardware, we applied 8-bit quantization to Llama-3.3-70B-Instruct. Smaller models were run without quantization.
Further, we use vLLM for more efficient memory management and higher throughput as this package enables sufficient batching of many requests at a time Kwon et al. (2023).
@@ Appendix C Evaluation
C.1.  Validation of Full-Text Persona Generation
We generate full-text personas based on the JSON-format persona descriptions with Gemini-2.5-flash. Although new, state-of-the-art LLMs are powerful and capable of dealing with large context windows, the generated full-text personas must be validated. LLMs are prone to hallucinations or misinterpreting the meaning of a response to a survey item when generating a full-text persona description. Although the maximum input token size of Gemini-2.5-flash exceeds 1 million tokens 333see Gemini-2.5-flash Documentation, more content in the context window might increase the likelihood of errors, hallucinations, or misinterpretation.
Apart from the minor misrepresentations and omissions found in the manually validated persona descriptions with up to 16 included attributes (TOP-2,-4,-8 and -16), we anecdotally observed that full-text personas with the largest sets of attributes (TOP-380) show major flaws. For multiple generated persona descriptions we observe generation loops, i.e., situations in which individual sentences in the persona description are repeated until the maximum token limit is reached. Therefore, we recommend using full-text personas with fewer attributes to avoid the use of these types of faulty personas.
C.2.  Hyperparameters for Random Forest Classifier
The Random Forest models utilized for classification tasks throughout all experiments were instantiated from the scikit-learn library’s RandomForestClassifier class. We selected the hyperparameters according to the findings of Miranda and Balbi (2025) who used the classifier in a similar task. To ensure consistency and comparability across experiments, the following hyperparameters were fixed: